OpenStudy (anonymous):

2x^2+7x-1=0 idk how to do this because i was sick for 2 days! :(

OpenStudy (anonymous):

I don't see a way to factor it. Try the Quadratic Formula

OpenStudy (anonymous):

when my techer waas doing it i was out of school...idk how to do it.... :/

OpenStudy (anonymous):

she never showed me, even tho i asked her to help me!

OpenStudy (anonymous):

\[x_{1,2}=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }=\frac{ 7\pm \sqrt{49+8} }{ 4 }=\frac{ 7\pm \sqrt{57} }{4 }\]